OpenStudy (anonymous):

can somebody help me to answer my assignment? substitution: integral dw over (4w^2-24w+27)^3\2

OpenStudy (freckles):

Did you try trig subsitution First you need to write the inside of that ( )^(3/2) has (f(w))^2+a

OpenStudy (freckles):

\[4w^2-24w+27=4(w^2-6w)+27=4(w^2-6w+9)+27-36=4(w-3)^2-9\]

OpenStudy (freckles):

\[=[2(w-3)]^2-3^2\]

OpenStudy (freckles):